Question 1 (20 marks) Gauteng Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Moulding Department is based on machine-hours, and the rate in the Painting Department is based on direct labour cost. At the beginning of the year, the company’s management made the following estimates: 6 PBA4807 MAY/JUNE 2019 PORTFOLIO EXAMINATION [TURN OVER] Department Moulding Painting Direct labour-hours 12,000 60,000 Machine-hours 70,000 8,000 Direct materials cost R510,000 R650,000 Direct labour cost 130,000 420,000 Manufacturing overhead cost 602,000 735,000 Job 205 was started on 1 August and completed on 10 August. The company’s cost records show the following information concerning the job: Department Moulding Painting Direct labour-hours 30 85 Machine-hours 110 20 Materials placed into production R470 R332 Direct labour cost 290 680 Required A. Compute the predetermined overhead rate used during the year in the Moulding Department. Compute the rate used in the Painting Department. B. Compute the total overhead cost applied to Job 205. C. What would be the total cost recorded for Job 205? If the job contained 50 units, what would be the cost per unit? D. At the end of the year, the records of Gauteng Potteryworks revealed the following actual cost and operating data for all jobs worked on during the year: Department Moulding Painting Direct labour-hours 10,000 62,000 Machine-hours 65,000 9,000 Direct materials cost R430,000 R680,000 Direct labour cost 108,000 436,000 Manufacturing overhead cost 570,000 750,000 7 PBA4807 MAY/JUNE 2019 PORTFOLIO EXAMINATION What was the amount of under- or overapplied overhead in each department at the end of the year?
Answer #1
A)Computation of Predetermined Overhead
Moulding Painting
Manufacturing overhead $602000 735000
Machine hours 70000 -
Direct labor cost $420000
Predetermined overhead rate    $8.6 $ 1.75
B. Computation of Total Overhead Cost Applied to job 205
Total overhead cost applied to job
Molding = 110 * 8.6 = $ 946
Painting = 680 * $1.75 = $ 1190
Total overhead applied = $2136
C. Computation of Total Cost recorded for the Job
Department Total
Molding Painting
direct material 470 332 802
direct labor 290 680 970
overhead applied 946 1190 2136
Total 1706 2202 3908
Unit Cot of Product, If 50 Unit in Job
unit product cost = 3908 / 50 =$78.16 per unit
4).
Molding Painting
Actual overhead Incurred 570000 750000
Overhead applied
65000 *8.6] 559000
[436000*1.75] 763000
overhead underapplied /(overapplied) 11000 13000
Underapplied Overapplied